the sample population is normally distributed, ¯x=42.5, σ = 3, and n = 30. What is the 90% confidence interval for µ?


1
Expert's answer
2022-04-11T15:01:31-0400

"90\%CI=(\bar x-z_{0.05}\frac{\sigma}{\sqrt{n}},\bar x+z_{0.05}\frac{\sigma}{\sqrt{n}})="

"=(42.5-1.645\frac{3}{\sqrt{30}},42.5+1.645\frac{3}{\sqrt{30}})=(41.60,43.40)."